# Topology optimization of nonlinear forced response curves via reduction on spectral submanifolds

### v1

@@ 5.2.1 Hardening optimization
Since one of the SN bifurcation points lies close to the peak of the FRC, the variation of the FRC peak with respect to the forcing amplitude $\epsilon$ can be studied via the SN bifurcation curves. As seen in the lower-right panel of Fig. 6, we allow for the variations in the forcing amplitude with $\epsilon=[0.01,1.2]\epsilon_{o}$ with $\epsilon_{o}=0.01$ to investigate how the SN bifurcations evolve with varying $\epsilon$. We can see that the variation amplitude of SN bifurcations of optimal layout $c_{\mathrm{nl}}$ is lower than that of $c_{\gamma}$. Consequently, as the forcing amplitude increases, the peak of FRC of $c_{\mathrm{nl}}$ remains lower than that of $c_{\gamma}$. To verify the convergence of the reduced-order model, we recomputed the SN bifurcation curves using an SSM reduction truncated at $\mathcal{O}(7)$. As shown in the same panel, the results confirm the convergence. Finally, the computational performance for $\gamma_{\mathrm{target}}=1\times 10^{-3}$ is summarized in Table 2.
@@
$c_{\gamma}$ $100\times 20$ 35 min 250 8.4 s/step
$c_{\text{nl}}$ $100\times 20$ 21 min 154 8.0 s/step
@@ 5.2.2 Softening Optimization
Similar to the case of hardening optimization, we perform parameter continuation of the SN bifurcation points with varying forcing amplitude $\epsilon=[0.01,1.2]\epsilon_{o}$. The projections of the obtained SN bifurcation curves are shown in the lower-right panel of Fig 7, from which we can see that the vibration amplitude of SN bifurcation points of $c_{\mathrm{nl}}$ is lower than that of $c_{\gamma}$. Therefore, the optimized structure $c_{\mathrm{nl}}$ demonstrates consistently enhanced resistance to external disturbances compared to $c_{\gamma}$ under various forcing amplitudes. Finally, we examine the computational performance for the case of $\gamma_{\mathrm{target}}=-1\times 10^{-3}$, as summarized in Table 3.
@@
$c_{\gamma}$ $100\times 20$ 29 min 203 8.5 s/step
$c_{\text{nl}}$ $100\times 20$ 20 min 154 7.9 s/step
@@ 5.3 Tuning saddle-node (SN) bifurcations on FRC
The final example is a microbeam in a mass sensor sun_new_2025 . The goal is to maximize the vibration amplitude of the microbeam to improve its measurement accuracy. Meanwhile, to prevent excessive amplitude fluctuations, we control the SN bifurcation points on the FRC to maximize the region without bistability. As seen in Fig. 8, the design domain of the beam has a length of 600 $\mathrm{\mu m}$ and a height of 100 $\mathrm{\mu m}$. A fixed region in the middle accounts for 20% of the design domain. The left end of the beam is fixed, while the right end is constrained in the $x$-direction. A periodic excitation is applied at the center of the right edge.
@@
Similar to the case of the MBB beam, a mesh of $120\times 20$ elements is used in the FE model, leading to a system with nn= 5019 DOF in (1). The damping ratio of initial layout $\xi$ = 0.1%, which leads to the values of damping constants $\alpha=1.97$ and $\beta=2.77\times 10^{-7}$. During the optimization iteration process, the forcing amplitude is given as $f^{\mathrm{ext}}=2\times 10^{8}\ \mathrm{ng\cdot\mu m/ms^{2}}$. In the following computation, we set $\epsilon=0.01$ unless otherwise stated. The target values for constraints in formulation (32) are set as $\omega_{X,\text{target}}=1500~\mathrm{kHz}$, and $A_{\mathrm{target}}=40\%$. Based on these parameter settings, we present the optimization result of optimization formulation (32) with different values of $b_{\mathrm{target}}$.
@@
Figure 8: Initial layout of the half microbeam considered in the example of Sec. 5.3. The total domain is 600 $\mathrm{\mu m}$ long and 100 $\mathrm{\mu m}$ high, and includes a non-design region of 100 $\mathrm{\mu m}$ in length. The gray area represents the designable region, while the black area indicates the fixed non-design region.
@@
We consider three cases $b_{\mathrm{target}}\in\{2,1,0.1\}$. The obtained results are shown in Fig. 9. From the first row, we observe that the optimal layouts evolve with decreasing $b_{\mathrm{target}}$, showing both changes in topology and clearer material boundaries. The second row shows the variation of $|b|$ during the iterative process. At the beginning of the iteration, the value of $|b|$ is 0, which indicates that the SN bifurcation point has not yet occurred. As the number of iteration steps increases, the vibration amplitude of the system becomes larger, the SN bifurcation points emerge, the value of $|b|$ is greater than 0, and it finally tends to the value of $b_{\mathrm{target}}$. Further, as the value of $b_{\mathrm{target}}$ decreases, more iterations are required for the convergence of $|b|$. Finally, we consider the FRCs of the three optimal layouts, as shown in the third row of Fig. 9. In these plots, the vertical axis $\|z_{\rm{out1}}\|_{\infty}$ denotes the vibration amplitude at the node located at $(x,y)=(600,50)$. We observe that two SN bifurcation points get closer as the value of $b_{\mathrm{target}}$ decreases.
@@
Figure 9: Optimization results of formulation (32) with different values of $b_{\mathrm{target}}$: 2 (left column), 1 (middle column), and 0.1 (right column). Top row: Optimal layouts. Middle row: Evolution of $|b|$ during the optimization iterations. Bottom row: FRCs of the optimal layouts computed using third-order SSM reduction.
@@
Now we present a close look at the iterative process for the case of $b_{\mathrm{target}}$ = 0.1. The evolution of the objective function $\rho_{\max}$ in optimization formulation (32) is shown in the left panel of Fig. 10. It can be observed that the value of the objective function increases during the initial iterations and gradually stabilizes as the optimization progresses. The corresponding evolution of the FRCs is shown in the right panel of Fig.10, where the peak amplitude of the FRCs also increases with the number of iterations. The computational performance of the above iterative process is summarized in Table 4.
@@
Figure 10: Evolution of the optimization process for $b_{\mathrm{target}}=0.1$. Left panel: convergence history of the objective function $\rho_{\max}$. Right panel: corresponding evolution of the FRCs over iterations.
@@
$120\times 20$ 58 min 301 11.6 s/step
Similar to the previous two examples, it is necessary to examine the convergence of the SSM-based model reduction. In Appendix D, we compute the FRC of the optimized structure shown in Fig. 9 using a seventh-order SSM-reduced model. The results indicate that the third-order SSM model has already achieved convergence.
@@ 6 Conclusion
In future work, we aim to extend this framework to the control of internal resonances, which often arise in nonlinear systems and significantly influence complex dynamic responses. In particular, internal resonances may induce Hopf bifurcations, which lead to quasi-periodic motions on invariant tori liang_bifurcation_2025 ; li_nonlinear_2022_part2 . We plan to explore how such modal interactions can be systematically controlled by manipulating reduced-order dynamics, including the regulation of Hopf bifurcations and the resulting quasi-periodic responses, to enable advanced control of multi-frequency behaviors in nonlinear structures.
@@ Appendix A The sensitivity of $\tilde{f}$
Appendix B The sensitivity of $b$
We first compute the sensitivity of $\rho_{\mathrm{SN}}$ and $\Omega_{\rm{SN}}$. Similar to compute $\rho_{\max}$, we compute the sensitivity of $\rho_{\rm{SN}}$ and $\Omega_{\rm{SN}}$ by taking derivative of (25), namely
$\rho_{\mathrm{SN}}^{\prime}=\frac{c_{1}}{c_{2}},\Omega_{\mathrm{SN}}^{\prime}=c_{3}-\frac{c_{4}}{c_{5}},$ (49)
@@ Appendix B The sensitivity of $b$
$c_{1}=$ $8\mathrm{Im}(\gamma)\mathrm{Im}(\gamma)^{\prime}\rho^{6}(\epsilon^{2}|\tilde{f}|^{2}-(\mathrm{Re}(\lambda)\rho+\mathrm{Re}(\gamma)\rho^{3})^{2})$ (50)
$+4\mathrm{Im}(\gamma)^{2}\rho^{6}(2\epsilon^{2}|\tilde{f}||\tilde{f}|^{\prime}-2(\mathrm{Re}(\lambda)\rho+\mathrm{Re}(\gamma)\rho^{3})(\mathrm{Re}(\lambda)^{\prime}\rho+\mathrm{Re}(\gamma)^{\prime}\rho^{3}))$
$-2(\epsilon^{2}|\tilde{f}|^{2}+2\mathrm{Re}(\lambda)\mathrm{Re}(\gamma)\rho^{4}+2\mathrm{Re}(\gamma)^{2}\rho^{6})$
$(2\epsilon^{2}|\tilde{f}||\tilde{f}|^{\prime}+2\mathrm{Re}(\lambda)^{\prime}\mathrm{Re}(\gamma)\rho^{4}+2\mathrm{Re}(\lambda)\mathrm{Re}(\gamma)^{\prime}\rho^{4}+4\mathrm{Re}(\gamma)\mathrm{Re}(\gamma)^{\prime}\rho^{6}),$
$c_{2}=$ $2(\epsilon^{2}|\tilde{f}|^{2}+2\mathrm{Re}(\lambda)\mathrm{Re}(\gamma)\rho^{4}+2\mathrm{Re}(\gamma)^{2}\rho^{4}+2\mathrm{Re}(\gamma)^{2}\rho^{6})$
@@
$+8\mathrm{Im}(\gamma)^{2}(\mathrm{Re}(\lambda)\rho+\mathrm{Re}(\gamma)\rho^{3})(\mathrm{Re}(\lambda)+3\mathrm{Re}(\gamma)\rho^{2})\rho^{6},$
$c_{3}=$ $\mathrm{Im}(\lambda)^{\prime}+\mathrm{Im}(\gamma)^{\prime}\rho^{2}+2\mathrm{Im}(\gamma)\rho\rho^{\prime},$
$c_{4}=$ $2\epsilon^{2}|\tilde{f}||\tilde{f}|^{\prime}-2(\mathrm{Re}(\lambda)\rho+\mathrm{Re}(\gamma)\rho^{3})$
$(\mathrm{Re}(\lambda)^{\prime}\rho+\mathrm{Re}(\lambda)\rho^{\prime}+\mathrm{Re}(\gamma)^{\prime}\rho^{3}+3\mathrm{Re}(\gamma)\rho^{2}\rho^{\prime})$
$-2(\mathrm{Im}(\lambda)-\Omega_{\mathrm{SN}}+\mathrm{Im}(\gamma)\rho^{2})^{2}\rho\rho^{\prime},$
$c_{5}=$ $2(\mathrm{Im}(\lambda)-\Omega+\mathrm{Im}(\gamma)\rho^{2})\rho^{2}.$
Then we consider the sensitivity of $\boldsymbol{A}_{\mathrm{SN}}$ and $\boldsymbol{B}_{\mathrm{SN}}$. By using the chain rule to the expressions in (29), we obtain
$\boldsymbol{A}_{\mathrm{SN}}^{\prime}$ $=\partial_{\boldsymbol{\mu}}\boldsymbol{A}_{\mathrm{SN}}+\partial_{\Omega_{\mathrm{SN}}}\boldsymbol{A}_{\mathrm{SN}}\Omega_{\mathrm{SN}}^{\prime}+\partial_{\rho_{\mathrm{SN}}}\boldsymbol{A}_{\mathrm{SN}}\rho_{\mathrm{SN}}^{\prime},$ (51)
$\boldsymbol{B}_{1}^{\prime}$ $=\partial_{\boldsymbol{\mu}}\boldsymbol{B}_{1}+\partial_{\Omega_{\mathrm{SN}}}\boldsymbol{B}_{1}\Omega_{\mathrm{SN}}^{\prime}+\partial_{\rho_{\mathrm{SN}}}\boldsymbol{B}_{1}\rho_{\mathrm{SN}}^{\prime},$
$\boldsymbol{B}_{2}^{\prime}$ $=\partial_{\boldsymbol{\mu}}\boldsymbol{B}_{2}+\partial_{\Omega_{\mathrm{SN}}}\boldsymbol{B}_{2}\Omega_{\mathrm{SN}}^{\prime}+\partial_{\rho_{\mathrm{SN}}}\boldsymbol{B}_{2}\rho_{\mathrm{SN}}^{\prime},$
where
$\partial_{\boldsymbol{\mu}}\boldsymbol{A}$ $=\begin{pmatrix}\mathrm{Re}(\lambda)^{\prime}+3\mathrm{Re}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{2}&-\left(\mathrm{Im}(\lambda)^{\prime}+\mathrm{Im}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{2}\right)\rho_{\mathrm{SN}}\\ 2\mathrm{Im}(\gamma)^{\prime}\rho_{\mathrm{SN}}+\left(\mathrm{Im}(\lambda)^{\prime}+\mathrm{Im}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{2}\right)/\rho_{\mathrm{SN}}&\mathrm{Re}(\lambda)^{\prime}+\mathrm{Re}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{2}\end{pmatrix}$, (52)
$\partial_{\Omega_{\mathrm{SN}}}\boldsymbol{A}$ $=\begin{pmatrix}0&\rho_{\mathrm{SN}}\\ -1/\rho_{\mathrm{SN}}&0\end{pmatrix}$,
$\partial_{\rho_{\mathrm{SN}}}\boldsymbol{A}$ $=\begin{pmatrix}6\mathrm{Re}(\gamma)\rho_{\mathrm{SN}}&-\mathrm{Im}(\lambda)+\Omega_{\mathrm{SN}}-3\mathrm{Im}\gamma\rho_{\mathrm{SN}}^{2}\\ 2\mathrm{Im}(\gamma)-(\mathrm{Im}(\lambda)-\Omega_{\mathrm{SN}})/\rho_{\mathrm{SN}}^{2}+\mathrm{Im}(\gamma)&2\mathrm{Re}(\gamma)\rho_{\mathrm{SN}}\end{pmatrix}$,
$\partial_{\boldsymbol{\mu}}\boldsymbol{B}_{1}$ $=\begin{pmatrix}6\mathrm{Re}(\gamma)^{\prime}\rho_{\mathrm{SN}}&0\\ 0&\mathrm{Re}(\lambda)^{\prime}\rho_{\mathrm{SN}}+\mathrm{Re}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{3}\end{pmatrix}$,
$\partial_{\Omega_{\mathrm{SN}}}\boldsymbol{B}_{1}$ $=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$,
$\partial_{\rho_{\mathrm{SN}}}\boldsymbol{B}_{1}$ $=\begin{pmatrix}6\mathrm{Re}(\gamma)&0\\ 0&\mathrm{Re}(\lambda)+3\mathrm{Re}(\gamma)\rho_{\mathrm{SN}}^{2}\end{pmatrix}$,
$\partial_{\boldsymbol{\mu}}\boldsymbol{B}_{2}$ $=\begin{pmatrix}2\mathrm{Im}(\gamma)^{\prime}-2(\mathrm{Im}(\lambda)^{\prime}+\mathrm{Im}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{2})/\rho_{\mathrm{SN}}^{2}&-(\mathrm{Re}(\lambda)^{\prime}+\mathrm{Re}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{2})/\rho_{\mathrm{SN}}\\ -(\mathrm{Re}(\lambda)^{\prime}+\mathrm{Re}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{2})/\rho_{\mathrm{SN}}&\mathrm{Im}(\lambda)^{\prime}+\mathrm{Im}(\gamma)^{\prime}\rho_{\mathrm{SN}}^{2}\end{pmatrix}$,
$\partial_{\Omega_{\mathrm{SN}}}\boldsymbol{B}_{2}$ $=\begin{pmatrix}2/\rho_{\mathrm{SN}}^{2}&0\\ 0&-1\end{pmatrix}$,
$\partial_{\rho_{\mathrm{SN}}}\boldsymbol{B}_{2}$ $=\begin{pmatrix}4(\mathrm{Im}(\lambda)-\Omega_{\mathrm{SN}})/\rho_{\mathrm{SN}}^{3}&\mathrm{Re}(\lambda)/\rho_{\mathrm{SN}}^{2}-\mathrm{Re}(\gamma)\\ \mathrm{Re}(\lambda)/\rho_{\mathrm{SN}}^{2}-\mathrm{Re}(\gamma)&2\mathrm{Im}(\gamma)\rho_{\mathrm{SN}}\end{pmatrix}$.
Thus, the sensitivity of b with respect to the design variable $\boldsymbol{\mu}$ is given by
$b^{\prime}=\boldsymbol{\psi}_{\mathrm{SN}}^{\prime}\boldsymbol{B}_{\mathrm{SN}}(\boldsymbol{\phi}_{\mathrm{SN}},\boldsymbol{\phi}_{\mathrm{SN}})+\boldsymbol{\psi}_{\mathrm{SN}}\boldsymbol{B}_{\mathrm{SN}}^{\prime}(\boldsymbol{\phi}_{\mathrm{SN}},\boldsymbol{\phi}_{\mathrm{SN}})+2\boldsymbol{B}_{\mathrm{SN}}(\boldsymbol{\phi}_{\mathrm{SN}},\boldsymbol{\phi}_{\mathrm{SN}}^{\prime}),$ (53)
where $\boldsymbol{\psi}^{\prime}_{\mathrm{SN}}$ and $\boldsymbol{\phi}^{\prime}_{\mathrm{SN}}$ are obtained by
$\begin{pmatrix}\boldsymbol{A}^{\mathrm{T}}&\boldsymbol{\psi}_{\mathrm{SN}}\\ \boldsymbol{\psi}_{\mathrm{SN}}^{\mathrm{T}}&0\end{pmatrix}\begin{pmatrix}\boldsymbol{\psi}^{\prime}_{\mathrm{SN}}\\ \xi\end{pmatrix}$ $=\begin{pmatrix}-\boldsymbol{A}^{\prime\mathrm{T}}\boldsymbol{\psi}_{\mathrm{SN}}\\ 0\end{pmatrix}$, (54)
$\begin{pmatrix}\boldsymbol{A}&\boldsymbol{\psi}_{\mathrm{SN}}\\ \boldsymbol{\psi}_{\mathrm{SN}}^{\mathrm{T}}&0\end{pmatrix}\begin{pmatrix}\boldsymbol{\phi}^{\prime}_{\mathrm{SN}}\\ \xi\end{pmatrix}$ $=\begin{pmatrix}-\boldsymbol{A}^{\prime}\boldsymbol{\phi}_{\mathrm{SN}}\\ -\boldsymbol{\phi}_{\mathrm{SN}}^{\mathrm{T}}\boldsymbol{\psi}_{\mathrm{SN}}^{\prime}\end{pmatrix}$.
Appendix C The convergence validation of SSM-based reduction in Sec. 5.1

### v2

@@ 5.2.1 Hardening optimization
Since one of the SN bifurcation points lies close to the peak of the FRC, the variation of the FRC peak with respect to the forcing amplitude $\epsilon$ can be studied via the SN bifurcation curves. As seen in the lower-right panel of Fig. 7, we allow for the variations in the forcing amplitude with $\epsilon=[0.01,1.2]\epsilon_{o}$ with $\epsilon_{o}=0.01$ to investigate how the SN bifurcations evolve with varying $\epsilon$. We can see that the variation amplitude of SN bifurcations of optimal layout $c_{\mathrm{nl}}$ is lower than that of $c_{\gamma}$. Consequently, as the forcing amplitude increases, the peak of FRC of $c_{\mathrm{nl}}$ remains lower than that of $c_{\gamma}$. To verify the convergence of the reduced-order model, we recomputed the SN bifurcation curves using an SSM reduction truncated at $\mathcal{O}(7)$. As shown in the same panel, the results confirm the convergence. Finally, the computational performance for $\gamma_{\mathrm{target}}=1\times 10^{-3}$ is summarized in Table 2Here the stopping criterion for the optimization of $c_{\text{nl}}$ is the same as that of Sec 5.1. As for the optimization problem with objective $c_{\gamma}$, since the objective function $c_{\gamma}\approx 0$ during optimization iterations, it is not suitable for judgment to optimize convergence. Thus, we take satisfaction of all constraints and the convergence of the design variable as the judgment, i.e., $\max|{\bar{\mathbf{\mu}}^{(i)}-\bar{\mathbf{\mu}}^{(i-1)}}|<10^{-2}$.
@@
$c_{\gamma}$ $100\times 20$ 35 mins 250 8.4 s/step
$c_{\text{nl}}$ $100\times 20$ 21 mins 154 8.0 s/step
@@ 5.2.2 Softening Optimization
Similar to the case of hardening optimization, we perform parameter continuation of the SN bifurcation points with varying forcing amplitude $\epsilon=[0.01,1.2]\epsilon_{o}$. The projections of the obtained SN bifurcation curves are shown in the lower-right panel of Fig 8, from which we can see that the vibration amplitude of SN bifurcation points of $c_{\mathrm{nl}}$ is lower than that of $c_{\gamma}$. Therefore, the optimized structure $c_{\mathrm{nl}}$ demonstrates consistently enhanced resistance to external disturbances compared to $c_{\gamma}$ under various forcing amplitudes. Finally, we examine the computational performance for the case of $\gamma_{\mathrm{target}}=-1\times 10^{-3}$, as summarized in Table 3The stopping criteria for $c_{\gamma}$ and $c_{\text{nl}}$ here are the same as that of Table 2.
@@
$c_{\gamma}$ $100\times 20$ 29 mins 203 8.5 s/step
$c_{\text{nl}}$ $100\times 20$ 20 mins 154 7.9 s/step
Similar to Sec. 5.1, we find that modes are well separated such that neither modal collisions nor internal resonances occur in this example. Detailed results and analysis are shown in Appendix D.1. In addition, results in the appendix show that the initial and optimized structures exhibit similar linear modal shapes for the first mode, which again indicates that the consistency of master mode is ensured in the optimization process.
The optimization results presented in this subsection show that the optimal layouts consistently exhibit hardening behavior when $\gamma_{\text{target}}>0$ and softening behavior when $\gamma_{\text{target}}<0$. This indicates that the proposed optimization framework can steer the nonlinear dynamic response of the structure toward a hardening/softening behavior. Further, a comparison with the reference formulation in Eq. (23) shows that Eq. (21) achieves concurrent minimization of the peak amplitude of FRC and the regulation of backbone curve.
@@ 5.3 Tuning saddle-node (SN) bifurcations on FRC
The final example is a microbeam in a mass sensor Sun et al. (2025). The goal is to maximize the vibration amplitude of the microbeam to improve its measurement accuracy. Meanwhile, to prevent excessive amplitude fluctuations, we control the SN bifurcation points on the FRC to maximize the region without bistability. As seen in Fig. 9, the design domain of the beam has a length of 600 $\mathrm{\mu m}$ and a height of 100 $\mathrm{\mu m}$. A fixed region in the middle accounts for 20% of the design domain. The left end of the beam is fixed, while the right end is constrained in the $x$-direction. A periodic excitation is applied at the center of the right edge.
@@
Similar to the case of the MBB beam, a mesh of $120\times 20$ elements is used in the FE model, leading to a system with nn= 5019 DOF in (1). The damping ratio of initial layout $\xi$ = 0.1%, which leads to the values of damping constants $\alpha=1.97$ and $\beta=2.77\times 10^{-7}$. During the optimization iteration process, the forcing amplitude is given as $f^{\mathrm{ext}}=2\times 10^{8}\ \mathrm{ng\cdot\mu m/ms^{2}}$. In the following computation, we set $\epsilon=0.01$. The projection parameter is initialized at $\sigma_{0}=10$ and multiplied by 1.5 every 60 iterations until it reaches a maximum value of $\sigma_{\max}=50$, while the parameter $p$ is initialized at $p_{0}=1$ and increased by one every 60 iterations, up to a maximum value of $p_{\max}=8$. The target values for constraints in formulation (29) are set as $\omega_{Y,\text{target}}=100~\mathrm{kHz}$, $\omega_{X,\text{target}}=500~\mathrm{kHz}$, and $A_{\mathrm{target}}=40\%$. Based on these parameter settings, we present the optimization result of optimization formulation (29) with different values of $d_{\mathrm{target}}$.
@@
Figure 9: Initial layout of the half microbeam considered in the example of Sec. 5.3. The total domain is 600 $\mathrm{\mu m}$ long and 100 $\mathrm{\mu m}$ high, and includes a non-design region of 20 $\mathrm{\mu m}$ in height located at the center of the domain. The gray area represents the designable region, while the black area indicates the fixed non-design region.
@@
We consider three cases $d_{\mathrm{target}}\in\{0.1,0.01,0\}$. The obtained results are shown in Fig. 10. From the first row, we observe that the optimal layouts evolve with decreasing $d_{\mathrm{target}}$, exhibiting noticeable changes in local shape features. The second row shows the variation of $|\Omega_{\text{SN1}}-\Omega_{\text{SN2}}|$ during the iterative process. At the beginning of the iteration, the value of $|\Omega_{\text{SN1}}-\Omega_{\text{SN2}}|$ is 0, which indicates that the SN bifurcation point has not yet occurred. As the number of iteration steps increases, the vibration amplitude of the system becomes larger, the SN bifurcation points emerge, the value of $|\Omega_{\text{SN1}}-\Omega_{\text{SN2}}|$ is greater than 0, and it finally tends to the value of $d_{\mathrm{target}}$. Further, when $d_{\mathrm{target}}=0$, more iterations are required for the convergence of $|\Omega_{\text{SN1}}-\Omega_{\text{SN2}}|$. The abrupt change of $|\Omega_{\text{SN1}}-\Omega_{\text{SN2}}|$ observed during the iteration process is due to the update of the penalization exponent $p$ and the projection parameter $\sigma$. Finally, we consider the FRCs of the three optimal layouts, as shown in the third row of Fig. 10. In these plots, the vertical axis $\|z_{\rm{out1}}\|_{\infty}$ denotes the vibration amplitude at the node located at $(x,y)=(600,50)$. We observe that two SN bifurcation points get closer as the value of $d_{\mathrm{target}}$ decreases.
@@
Figure 10: Optimization results of formulation (29) with different values of $d_{\mathrm{target}}$: 0.1 (left column), 0.01 (middle column), and 0 (right column). Top row: Optimal layouts. Middle row: Evolution of $|\Omega_{\text{SN1}}-\Omega_{\text{SN2}}|$ during the optimization iterations. Bottom row: FRCs of the optimal layouts computed using third-order SSM reduction.
@@
Now we present a close look at the iterative process for the case of $d_{\mathrm{target}}$ = 0. The evolution of the objective function $\rho_{\max}$ in optimization formulation (29) is shown in the left panel of Fig. 11. It can be observed that the value of the objective function increases during the initial iterations and gradually stabilizes as the optimization progresses. The corresponding evolution of the FRCs is shown in the right panel of Fig.11, where the peak amplitude of the FRCs also increases with the number of iterations. The computational performance of the above iterative process is summarized in Table 4Here, the stopping criterion for the optimization of $c_{\text{SN}}$ is similar to that of $c_{\text{nl}}$. Specifically, the stopping criterion is defined by the simultaneous satisfaction of all constraints and the steady convergence of the objective function $c_{\text{SN}}$.
@@
Figure 11: Evolution of the optimization process for $d_{\mathrm{target}}$ = 0. Left panel: convergence history of the objective function $\rho_{\max}$. Right panel: corresponding evolution of the FRCs over iterations.
@@
$120\times 20$ 53 mins 458 7.0 s/step
Similar to the previous two examples, it is necessary to examine the convergence of the SSM-based model reduction. In Appendix E.1, we compute the FRC of the optimized structure shown in Fig. 10 using a seventh-order SSM-reduced model. The results indicate that the third-order SSM model has already achieved convergence.
To justify the assumption of no internal resonance, we present in Appendix E.2 the evolution of the first three eigenfrequencies throughout the iterative process for the case of $d_{\mathrm{target}}=0$. From these results, we observe that modes are well separated and neither modal collisions nor internal resonance occur. In addition, the consistency of master mode during optimization is again validated in Appendix E.2.
The optimization results presented in this subsection indicate that decreasing the prescribed value of $d_{\mathrm{target}}$ in the optimization formulation leads to a reduced frequency distance between the two SN bifurcations on the FRC. This observation demonstrates the effectiveness of the SSM-based framework in regulating the length of segment of unstable periodic orbits on FRC.
@@ 6 Conclusion
In future work, we aim to extend this framework to the control of internal resonances, which often arise in nonlinear systems and significantly influence complex dynamic responses. In particular, internal resonances may induce Hopf bifurcations, which lead to quasi-periodic motions on invariant tori Liang et al. (2025); Li and Haller (2022). We plan to explore how such modal interactions can be systematically controlled by manipulating reduced-order dynamics, including the regulation of Hopf bifurcations and the resulting quasi-periodic responses, to enable advanced control of multi-frequency behaviors in nonlinear structures. The proposed framework can also be extended to realistic MEMS devices by incorporating anchors, electrodes, and actuation gaps, implemented through non-design regions, fixed boundary conditions, and appropriate optimization constraints. We leave this extension to future study.
@@ Appendix A The sensitivity of $\tilde{f}$
Appendix B The sensitivity of $\Omega_{\text{SN}}$
We first compute the sensitivity of $\rho_{\mathrm{SN}}$ by taking derivative of (26), namely
$\rho_{\mathrm{SN}}^{\prime}=\frac{c_{1}}{c_{2}},$ (47)
@@ Appendix B The sensitivity of $\Omega_{\text{SN}}$
$c_{1}=$ $8\mathrm{Im}(\gamma)\mathrm{Im}(\gamma)^{\prime}\rho_{\text{SN}}^{6}(\epsilon^{2}|\tilde{f}|^{2}-(\mathrm{Re}(\lambda)\rho_{\text{SN}}+\mathrm{Re}(\gamma)\rho_{\text{SN}}^{3})^{2})$ (48)
$+4\mathrm{Im}(\gamma)^{2}\rho_{\text{SN}}^{6}(2\epsilon^{2}|\tilde{f}||\tilde{f}|^{\prime}-2(\mathrm{Re}(\lambda)\rho_{\text{SN}}+\mathrm{Re}(\gamma)\rho_{\text{SN}}^{3})(\mathrm{Re}(\lambda)^{\prime}\rho_{\text{SN}}+\mathrm{Re}(\gamma)^{\prime}\rho_{\text{SN}}^{3}))$
$-2(\epsilon^{2}|\tilde{f}|^{2}+2\mathrm{Re}(\lambda)\mathrm{Re}(\gamma)\rho_{\text{SN}}^{4}+2\mathrm{Re}(\gamma)^{2}\rho_{\text{SN}}^{6})$
$(2\epsilon^{2}|\tilde{f}||\tilde{f}|^{\prime}+2\mathrm{Re}(\lambda)^{\prime}\mathrm{Re}(\gamma)\rho_{\text{SN}}^{4}+2\mathrm{Re}(\lambda)\mathrm{Re}(\gamma)^{\prime}\rho_{\text{SN}}^{4}+4\mathrm{Re}(\gamma)\mathrm{Re}(\gamma)^{\prime}\rho_{\text{SN}}^{6}),$
$c_{2}=$ $2(\epsilon^{2}|\tilde{f}|^{2}+2\mathrm{Re}(\lambda)\mathrm{Re}(\gamma)\rho_{\text{SN}}^{4}+2\mathrm{Re}(\gamma)^{2}\rho_{\text{SN}}^{4}+2\mathrm{Re}(\gamma)^{2}\rho_{\text{SN}}^{6})$
@@
$+8\mathrm{Im}(\gamma)^{2}(\mathrm{Re}(\lambda)\rho_{\text{SN}}+\mathrm{Re}(\gamma)\rho_{\text{SN}}^{3})(\mathrm{Re}(\lambda)+3\mathrm{Re}(\gamma)\rho_{\text{SN}}^{2})\rho_{\text{SN}}^{6}.$
Then, the sensitivity of $\Omega_{\rm{SN}}$ can be obtained by taking derivative of (18), namely
$\Omega_{\mathrm{SN}}^{\prime}=c_{3}-\frac{c_{4}}{c_{5}},$ (49)
where $c_{3},c_{4}$ and $c_{5}$ have been defined in (35).
Appendix C Supplementary materials for the example in Sec. 5.1
C.1 The convergence validation of SSM-based reduction